OpenStudy (anonymous):

A gallon if water is added to 6 quarts of a solution that is 50% acid. What percent of the new solution is acid?

OpenStudy (anonymous):

The old solution contains 6(0.5) = 3 quarts of acid. The gallon of water has no acid. Total amount of acid in new solution = 3 quarts. The old solution has volume 6 quarts. The gallon of water has volume 4 quarts. Total volume of new solution = 4 + 6 = 10 quarts. The new solution is 3/10 = 0.3 = 30% acid.